OpenStudy (anonymous):

Which of the following shows the step necessary to solve the equation x plus five and five-sixths equals six and one-half and its correct solution? Subtract six and one-half from both sides; two-thirds Add six and one-half to both sides; twelve and one-third Subtract five and five-sixths from both sides; two-thirds Add five and five-sixths to both sides; twelve and one-third

OpenStudy (anonymous):

help pleas

OpenStudy (alyssa_xo):

get the fractions into common denominators and isolate the variable

OpenStudy (anonymous):

\[x + 5 \frac{5}{6} = 6 \frac{1}{2}\] Subtract 5 5/6 from both sides to get x alone. \[x = 6 \frac{1}{2}-5\frac{5}{6}\] 1/2 is really the same thing as 3/6. I got this by multplying the top and bottom by 3. You can do this as long as you do it to the top and bottom. Now we have: \[x = 6 \frac{3}{6}-5\frac{5}{6}\] Take one from the 6 and add it to the 3/6 as 6/6. We can do this because 6/6 is the same thing as 1. \[x = 5 \frac{9}{6}-5\frac{5}{6}\] \[x = \frac{4}{6}\] Reduce \[x = \frac{2}{3}\] So...your first option. "Subtract six and one-half from both sides; two-thirds"

OpenStudy (anonymous):

No, sorry, my bad. The third option. " Subtract five and five-sixths from both sides; two-thirds" becasue that's the first step i did.